### v1

# Evaluating Compiler Optimization Impacts on zkVM Performance

Abstract.
Zero-knowledge proofs (ZKPs) are the cornerstone of programmable cryptography. They enable (1) privacy-preserving and verifiable computation across blockchains, and (2) an expanding range of off-chain applications such as credential schemes. Zero-knowledge virtual machines (zkVMs) lower the barrier by turning ZKPs into a drop-in backend for standard compilation pipelines. This lets developers write proof-generating programs in conventional languages (e.g., Rust or C++) instead of hand-crafting arithmetic circuits. However, these VMs inherit compiler infrastructures tuned for traditional architectures rather than for proof systems. In particular, standard compiler optimizations assume features that are absent in zkVMs, including cache locality, branch prediction, or instruction-level parallelism. Therefore, their impact on proof generation is questionable.
We present the first systematic study of the impact of compiler optimizations on zkVMs. We evaluate 64 LLVM passes, six standard optimization levels, and an unoptimized baseline across 58 benchmarks on two RISC-V–based zkVMs (RISC Zero and SP1). While standard LLVM optimization levels do improve zkVM performance (over 40%), their impact is far smaller than on traditional CPUs, since their decisions rely on hardware features rather than proof constraints. Guided by a fine-grained pass-level analysis, we slightly refine a small set of LLVM passes to be zkVM-aware, improving zkVM execution time by up to 45% (average +4.6% on RISC Zero, +1% on SP1) and achieving consistent proving-time gains. Our work highlights the potential of compiler-level optimizations for zkVM performance and opens new direction for zkVM-specific passes, backends, and superoptimizers.
1. Introduction
Zero-knowledge proofs (ZKPs) (Goldwasser et al., 1985) let a prover convince a verifier that a statement holds, optionally, without revealing anything beyond its truth. After decades as a theoretical concept, advances in proof systems have made ZKPs practical at scale (Parno et al., 2013; Groth, 2016; Gabizon et al., 2019). Still, two obstacles have thwarted adoption: (1) developers had to handcraft low-level arithmetic circuits (Bellés-Muñoz et al., 2023), which is both hard and error-prone (Chaliasos et al., 2024a), and (2) proving remains costly despite recent progress (Ernstberger et al., 2024a).
Zero-knowledge virtual machines (zkVMs) (Bruestle and Gafni, 2023; Labs, 2024b; Arun et al., 2023; Goldberg et al., 2021) address both issues by enabling developers to reuse well-established languages and toolchains. This is achieved by compiling ordinary programs (e.g., Rust) to a standard ISA such as RISC-V (Waterman et al., 2016), then emulating the binary to record an execution trace that a zkVM back-end proves. This design leverages tools such as LLVM (Lattner and Adve, 2004), avoids manually-written circuits, and benefits from techniques such as GPU proving (Ma et al., 2023; Succinct Labs, 2025), recursion (Kothapalli et al., 2022; Zero, 2025a), and precompiles for heavy primitives. zkVMs are already deployed in key blockchain applications (Labs, 2024c; Mutmainnah, 2024), and are increasingly adopted in off-chain systems, such as anonymous credentials and client-side proving (Ames et al., 2023). A prime example is Ethereum’s long-term roadmap which is heavily centered around zkVMs. Current proposals include replacing the EVM with RISC-V to accelerate proving (Buterin, 2025), advancing formal-verification efforts to harden zkVM correctness (Verified zkEVM Initiative, 2025), and initiatives such as ETHProofs that monitor progress across both fronts (Ethereum Foundation, 2025).
Despite this progress, performance remains the central hurdle and a competitive factor among zkVM vendors (Guibas, 2024; Ethereum Foundation, 2025). zkVMs’ architecture adopts a cost model fundamentally different from traditional CPUs, where every instruction becomes a set of constraints to be proved. Yet, compilers still optimize for hardware features that are not relevant to zkVMs, such as caches, out-of-order execution, and instruction-level parallelism. All these raise the following question: How do compiler-level techniques (i.e., optimizations) influence zkVM performance, and to what extent?
This open question motivates this study, which aims to investigate how traditional optimizations influence zkVM performance. To do so, we seek answers for the following three research questions:
  1. RQ1
    How do standard compiler optimizations impact zkVM performance when applied individually? (Section 4.1)
  2. RQ2
    Are there combinations of optimizations that yield significant performance improvements or degradations? (Section 4.2)
  3. RQ3
    How does the impact of optimizations on zkVMs compare to traditional architectures? (Section 4.3)
[…]
Contributions. We make the following contributions:
  • We present the first systematic study of how traditional compiler optimizations (both individual passes and their combinations) affect zkVM performance, and compare the results to traditional CPUs.
  • We analyze selected optimizations with notable performance effects on zkVMs, extract insights about the underlying root causes, and use them to refine these optimizations for better zkVM performance.
  • We enumerate the implications of our findings, and discuss potential future directions on improving zkVM performance via compiler-level techniques.
Summary of findings. Some of our representative findings are: (1) default optimization levels (e.g., -O3) improve both zkVM execution and proving time by over 40%; (2) however, these gains are significantly smaller than on traditional CPUs, as many individual passes exhibit opposing effects, mainly due to wrong decision heuristics for inlining, loop unrolling, branch elimination, or mismatches in the cost models for arithmetic and memory operations; and (3) autotuning proves to be an effective approach for performance-critical zkVM programs (up to $2.2\times$ speedup compared to -O3).
Implications. Building on our findings, we identify root causes of inefficiencies in certain passes and implement slight modifications to an existing compiler toolchain (i.e., LLVM). The -O3 level in our modified LLVM outperforms the original LLVM’s -O3 in most benchmarks, achieving up to 45% faster zkVM execution (avg. +4% on RISC Zero; +1% on SP1) while also delivering consistent improvements in proving time.
2. Background
Zero-Knowledge Proofs and zkVMs. A ZKP (Goldwasser et al., 1985) is a cryptographic primitive that allows a prover to convince a verifier that a statement is valid, optionally, without revealing anything beyond its truth. A ZKP provides completeness, soundness, and optionally zero knowledge. In the past decade, adoption of ZKPs have accelerated with SNARKs (Parno et al., 2013; Groth, 2016), which provide non-interactive, succinct proofs with fast verification and practical proving costs. ZKPs have enabled many novel applications (Ernstberger et al., 2024b; Lavin et al., 2024) such as private transactions (Ben Sasson et al., 2014) and credential systems (Baldimtsi et al., 2024; Frigo and Shelat, 2024).
SNARKs target circuit satisfiability (Parno et al., 2013), representing statements as arithmetic circuits. Early ZK development used low-level ZK DSLs (Bellés-Muñoz et al., 2023), where a compiler arithmetizes the circuit into a concrete constraint system (e.g., PLONK-style (Gabizon et al., 2019)) and produces a witness generator. Then, the prover uses the constraints and witness to produce a proof. Writing low-level ZK circuits is not only error-prone (Chaliasos et al., 2024a), but it also requires significant expertise. Zero-knowledge virtual machines (zkVMs) (Ben-Sasson et al., 2014) raise the abstraction by compiling source code (e.g., Rust) to a target ISA (often RISC-V (Waterman et al., 2016)(Labs, 2024b; Bruestle and Gafni, 2023). Then, an emulator executes the binary and records an execution trace in the proper arithmetization, and the back-end produces a proof. In this context, developers use well-established compilation toolchains (e.g., LLVM), and the zkVM exposes precompiles for heavy primitives (e.g., SHA-2) to replace thousands of generic instructions with optimized circuits. To make zkVMs practical for computation-heavy applications (e.g., ZK-Rollups (Barry Whitehat, 2018)), much of the literature focuses on improving both execution and proof generation (Arun et al., 2023; Setty et al., 2023; Goldberg et al., 2021; Labs, 2024b; Bruestle and Gafni, 2023). Figure 1 contrasts native execution, ZK DSLs, and zkVMs. For an in-depth introduction of ZKPs, we refer the reader to (Thaler et al., 2022).
Figure 1. Architecture comparison between native execution, ZK DSLs, and zkVMs.
zkVMs vs. traditional architectures. When generating the execution trace and proving it on a zkVM, the cost model differs significantly from traditional CPU architectures, such as x86 or ARM. Most instructions have near-uniform cost, and there is no instruction latency due to cache hits/misses and other micro-architectural effects. Memory access is relatively cheaper on zkVMs, but page-ins and page-outs can be significantly more expensive. Further, there is no parallelism or multithreading, and floating-point operations are not supported natively, so they need to be emulated, which makes them extremely costly. We detail a full list of differences in Appendix A. Notably, zkVM execution and proving are orders of magnitude slower than native execution (milliseconds vs. seconds to hours, cf. Figure 14 Appendix A).
Compiler optimizations and zkVMs. Compiler optimizations are semantics-preserving transformations that aim to improve speed, reduce size, or enhance other attributes of code (Torczon and Cooper, 2007; Aho et al., 2006). We present two motivating examples where optimizations that aid conventional CPUs degrade zkVM performance, exposing a mismatch between hardware-oriented assumptions and the zkVMs’ proof-centric cost model.
Strength reduction. LLVM’s strength reduction replaces division instruction with a shift-and-add sequence (Figure 2(a)). On conventional CPUs, this is beneficial because division instruction is often costlier than the equivalent bitwise instructions (Fog, 2022). Our x86 measurements show that the transformed code is $3.5\times$ faster. On zkVMs, the effect reverses, notably proving is 40% slower in RISC Zero (Bruestle and Gafni, 2023). The reason is the cost model mismatch: on zkVMs, every operation contributes to the number of constraints in the proof, and all instructions have roughly uniform cost. Replacing a single division with multiple bitwise operations therefore increases the constraint count, making proving slower.
// unoptimized version
example::div:
        li t0, 8
        div a0, a0, t0
        ret
// optimized version
example::div:
        srai a1, a0, 31
        srli a1, a1, 29
        add a0, a0, a1
        srai a0, a0, 3
        ret
(a) CPU-oriented instruction selection can hurt zkVMs: on RISC Zero, the ‘optimized’ form takes 8 cycles vs. 4 unoptimized
/* unoptimized version */
int i, a[N], b[N];
for (i = 0; i < N; i++) {
    a[i] = 1;
    b[i] = 2;
/* optimized version */
int i, a[N], b[N];
for (i = 0; i < N; i++) {
    a[i] = 1;
for (i = 0; i < N; i++) {
    b[i] = 2;
(b) Loop fission often helps CPUs but can hurt zkVMs by duplicating loop-control work. NN is set to 10485761048576.
Figure 2. Optimizations with negative effects on zkVMs.
Loop fission is an optimization that splits a loop into multiple ones to enhance locality and reduce cache misses (Figure 2(b)). On x86, the transformed code is $\sim$8% faster. On zkVMs, the effect flips: duplicating loop control increases proof constraints, raising proving time by 5% on SP1 (Labs, 2024b). This underscores that cache-centric assumptions do not transfer to zkVMs’ model.
3. Methodology
Selection of zkVMs. We evaluate two zkVMs targeting RISC-V (Waterman et al., 2016) (SP1 (Labs, 2024b) and RISC Zero (Bruestle and Gafni, 2023)), chosen for maturity and widespread use in production. We also considered zkVMs with non-standard ZK-oriented ISAs such as Cairo (Goldberg et al., 2021) and Valida (Thomas et al., 2025), but these stacks bypass existing compiler infrastructures and rely on less stable transpilers.
The selected zkVMs can directly run C and Rust programs that are lowered to RISC-V via an existing compiler toolchain, in particular LLVM (Lattner and Adve, 2004). Therefore, our work specifically investigates the impact of traditional compiler optimizations implemented in LLVM on zkVMs.
3.1. Metrics
We evaluate the impact of optimizations on zkVMs using three zkVM metrics: cycle count, zkVM execution time, and proving time. Cycle count is the total cycles of the program run on the zkVM (also known as the guest program), and is analogous to CPU cycles. zkVM execution time is the wall-clock time for the executor to replay the guest and produce the full execution trace. Proving time is the wall-clock time for the prover to generate a proof and is typically larger than execution time. For native runs (RQ3), we measure native execution time to contrast optimization effects on zkVMs and traditional architectures.
3.2. Benchmark Selection
We combine standard benchmarks along with zkVM-focused workloads to cover both large, optimization-sensitive programs and common zkVM payloads. In total, our benchmark suite consists of 58 programs written in Rust and C/C++. We use PolyBench (Pouchet, 2010; Ferrer, 2020), NPB (Division, 2024; Martins et al., 2025), and SPEC CPU 2017 (Bucek et al., 2018), which include compute and memory-intensive programs. From industry benchmark suites, we include the a16z crypto zkVM, the Succinct Labs, and the RSP benchmarks (a16z crypto, 2024; Labs, 2024d, a). These capture typical cryptography-heavy workloads. We further add targeted programs, such regex-match, sha256, and loop-sum, to exercise loop-heavy, memory-heavy, and math/crypto patterns frequently seen in zkVMs. For the complete list of the benchmarks see Appendix B.
3.3. Analyzing Optimizations
To measure the impact of individual optimizations as well as standard optimization levels, we evaluate a total of 71 optimization profiles, which consist of 64 individual LLVM optimization passes, 1 unoptimized baseline profile, and 6 preset optimization levels (-O0, -O1, -Oz, etc.). For all 64 individual optimization passes and the baseline profile, we also disable Rust’s MIR optimizations by setting mir-opt-level to 0. The baseline profile has no optimizations applied and serves as a reference point for other optimization profiles.
To answer the RQs outlined in Section 1, we first measure each pass in isolation on RISC Zero and SP1 across our benchmark suite, collecting cycle count, executor time, and proving time (Section 4.1) and quantify the performance losses/gains compared to the baseline. Next, we explore optimization combinations via genetic autotuning with OpenTuner (Ansel et al., 2014) (Section 4.2). Then, for RQ3, we run the same optimization profiles on traditional CPU (i.e., x86) and compare the speedups with those obtained from zkVMs to pinpoint divergences caused by the different cost models (Section 4.3).
Setup. x86 and zkVM execution experiments run on an AMD EPYC 7742 with 500GB of RAM on Ubuntu 22.04.5. Proving runs on an AMD EPYC 7542 with 32GB RAM on Ubuntu 24.04.2, equipped with an NVIDIA GeForce RTX 4090 (24GB VRAM). We use SP1 4.1.4 (moongate-server 4.1.0), RISC Zero 1.2.4, Clang 19.1.7, LLVM 19.1.7, and Rust 1.85.0. For both zkVMs we used their default STARK-based proof systems.
3.4. Threats to Validity
Internal validity. Proving time on SP1 exhibits higher variance than on RISC Zero, partly due to the closed-source prover accessed via RPC and the associated network overhead. To mitigate this, We use at least 10 samples per configuration and increase the sample size for noisy cases. We aim to study broad optimization patterns rather than absolute performance. Cycle count is strongly but not perfectly correlated with proving time, so in edge cases it can over- or underestimate actual costs.
External validity. Our proving time measurements are accelerated on an NVIDIA RTX 4090. However, different GPUs or CPUs may change effects. We study two popular RISC-V zkVMs, but other zkVMs or designs (e.g., different ISAs) may behave differently. Extending the study to additional zkVMs is left for future work. Our RQ3 comparison is limited to x86. While ARM or RISC-V CPUs could provide additional insights, we choose x86 since it is currently the most widely used architecture for running zkVM workloads. Finally, results may not generalize to all workloads, but we mitigate this by selecting benchmarks spanning scientific, loop-intensive, and cryptographic domains.
[…]
4. Results
Figure 3. The impact of top 25 individual LLVM passes across all benchmarks, zkVMs (RISC Zero, SP1), and metrics (zkVM execution time, proving time, cycle count). Lines represent average impact; shaded areas show standard deviation. Higher values indicate better performance for proving time and zkVM execution time, while lower is better for cycle count.
In this section, we present the results for each research question outlined in Section 1.
4.1. RQ1: Impact of Individual LLVM Passes
Figure 3 presents the top 25 LLVM passes with the highest average impact (either positive or negative) across all benchmarks, zkVMs, and metrics. Each result reflects the performance difference relative to an unoptimized baseline with no passes applied. The remaining 39 passes have minimal impact and are omitted for brevity.
[…]
Inlining is beneficial for both zkVMs: the inline and always-inline passes lead to performance gains of 28.4% and 5.7% on RISC Zero, and 19.3% and 5.3% on SP1, respectively. Other passes, such as instcombine or sroa, perform well on some benchmarks, while degrade performance on others. Both zkVMs exhibit similar trends, with no pass consistently having opposite effects on RISC Zero and SP1.
Proving time. Proving time closely follows the trends observed in zkVM execution. The most beneficial pass is again inline (22.4% performance gain on both SP1 and RISC Zero). The most harmful pass remains licm, which increases proving time by 8.4% on SP1 and 13.5% on RISC Zero.
Cycle count. Similar trends are observed when examining cycle count. Inlining reduces cycle count with a 30.8% reduction on SP1 and a 29.9% on RISC Zero. The pass that increases the cycle count the most is licm, which explains its negative effect on zkVM execution and proving.
We compare cycle count with execution and proving times and find a strong linear correlation between them. On both zkVMs, the Pearson and Spearman correlation coefficients exceed 0.98. This confirms that increases in cycle count translate directly into longer execution and proving times. zkVM execution time and proving time show a similar strong linear correlation, with Pearson and Spearman scores above 0.98.
Figure 4. Counting the cases where each optimization pass leads to (1) severe performance gain ($\geq 5$%), moderate performance gain (between $2\%$ and $5\%$), moderate performance loss (between -2% and -5%), and severe performance loss ($\leq-5$%).
Table 1. Number of instances where optimized code leads to performance gains or losses in execution and proving time.
[…]
RISC Zero 370 437 302 241
SP1 314 124 347 174
Categories of effects. For completeness, Figure 4 shows the number of benchmarks where each optimization pass causes: (1) moderate performance loss (-5% to -2%), (2) severe loss ($\leq-5\%$), (3) moderate gain (2% to 5%), or (4) severe gain ($\geq 5\%$). Inlining consistently improves both execution and proving times, though a few benchmarks still see execution slowdowns (more details in Section 5). In contrast, loop optimizations (e.g., loop-extract, licm, loop-deletion, loop-unroll) often harm zkVM performance, especially on RISC Zero. This is because loop optimizations in LLVM operate in Loop Closed SSA (LCSSA) form, where values defined in a loop but used outside it are routed through phi nodes at the loop exit. This transformation often requires recomputing addresses (e.g., via getelementptr) or adding load/store instructions, especially when iterating over arrays, leading to extra memory operations that harm zkVM performance (Section 5). Finally, some passes, such as instcombine on RISC Zero execution, show a balanced impact, with similar numbers of benchmarks improving and regressing. This suggests that certain rewrites in instcombine may hurt zkVM performance and should be adjusted to better align with zkVMs’ model (Figure 2(a)).
Differences in zkVMs. Differences between zkVMs are evident in Table 1, which shows how often optimized code results in performance gains or losses in execution and proving time for each zkVM. Overall, individual optimizations appear to have a more detrimental effect on RISC Zero, especially in execution, where degradation occurs roughly $4\times$ more often compared to SP1. This could stem from differences in IRs and constraint-level optimizations across zkVM.
4.2. RQ2: Combinations of Optimizations
Standard optimization levels. Figure 5 shows the impact of standard optimization levels (-Ox flags) on execution and proving time across both zkVMs. For brevity, the effects on cycle count are omitted. Excluding -O0, no -Ox level negatively impacts performance relative to the baseline. In contrast, -O0 (i.e., Rust’s MIR optimizations) causes regressions in 19 programs on RISC Zero and 9 on SP1. On average, standard optimization levels improve execution time by 60.5% on RISC Zero and 47.3% on SP1, and proving time by 55.5% and 51.1%, respectively. The -O3 level consistently achieves the highest average performance gains, while -Oz (reducing binary size) shows the lowest.
Figure 5. Impact of standard optimization levels for zkVMs.
Autotuning. Overall, the existing -Ox flags have a positive effect on both zkVM execution and proving time. However, can we identify sequence of passes that outperform them? To explore this, we employ OpenTuner (Ansel et al., 2014), a genetic-based tuning framework that finds an optimal combination of compilation flags guided by a specific fitness function. Although zkVM execution and proving time are the ultimate metrics of interest, they are too expensive to evaluate repeatedly during autotuning. Instead, we use cycle count as a proxy, based on its strong linear correlation with both execution and proving established in RQ1 (Section 4.1). This choice offers two key advantages: (1) cycle count avoids measurement noise, and (2) it is fast to compute ($\sim$1 second on average).
[…]
(a) NPB benchmark suite
(b) Crypto benchmark suite
Figure 6. The speedup over -O3 (with LTO) achieved by tuning optimization passes for each program in the NPB and Crypto benchmark suites across both zkVMs.
However, we also observed more interesting behaviors: inline appears in 122 of the worst-performing sequences, often when followed by licm (23 out of 122). In contrast, the combination where licm follows inline occurs 50 times in best-performing sequences. Further analysis in Section 5 explores the root causes behind these context-sensitive effects.
[…]
Figure 6 shows the performance improvement of each program in the selected benchmark suites relative to -O3 with link-time optimizations (LTO). For the NPB benchmark suite (Figure 6(a)), the autotuned configurations achieve an average performance increase by 17% on SP1 in terms of both zkVM execution time and proving time. On RISC Zero, autotuning improves by 19% the execution time, and by 17% the proving time. The npb-sp program stands out, as the result of the OpenTuner achieves over $2\times$ speedup on both zkVMs for both metrics.
Moving to Figure 6(b), autotuning yields smaller, yet meaningful, gains for the cryptography benchmarks. On average, execution improves by 11.6% on RISC Zero and 10.4% on SP1, while proving time improves by 3.5% and 6.8%, respectively. The smaller gains are expected, as both zkVMs are already heavily optimized for cryptographic workloads through the use of precompiles. These are specialized, built-in circuits that implement certain expensive operations (e.g., elliptic-curve signature verification or hashing) directly in hardware-equivalent logic, instead of executing them step-by-step as regular instructions. This allows these operations to be proven far more efficiently, often in constant time.
Nevertheless, even for benchmarks that make use of such precompiles, such as ecdsa-verify, eddsa-verify, and keccak256, autotuning can still deliver performance improvements that exceed 10%.
The findings confirm the viability of autotuning on zkVMs, which has practical implications for optimizing performance-critical programs on zkVMs (see Section 6.2).
Security-critical bug in SP1. As a by-product, our autotuning efforts uncover a security-critical bug in SP1. In particular, while autotuning a specific program from the Crypto benchmark suite, we run into a certain sequence of passes that leads to 59% reduction in cycle count compared to -O3 (w/ LTO)! This, however, turns out to be a bug in SP1 that causes the zkVM to silently abort execution in mid-run while still producing a proof that passes verification. Such silent failures represent a serious vulnerability, as they allow incorrect program results to be proven as correct. We privately report this silent-failure case to the SP1 development team, who confirm it as a bug and then patch it.
4.3. RQ3: Comparison to Traditional CPUs
Figure 7. The average impact of each optimization on zkVM and x86 performance, measured across all benchmarks. Passes with an average effect below 2% are omitted.
We compare the impact of our studied LLVM passes and flags on zkVMs versus a traditional architecture, that is, x86.
Figure 7 shows the average impact of each optimization pass on zkVM and x86 performance, measured across all benchmarks. For brevity, we exclude passes with an average impact on zkVM execution time, proving time, and x86 execution time that is below 2%. For the majority of passes, the direction of their impact (positive or negative) is similar on both zkVMs and x86. However, the magnitude of the effect is significantly larger on x86. This could be explained by the fact that many LLVM optimizations fail to show their full potential on zkVMs. This is mainly due to a mismatch in the cost model or the underlying heuristics of LLVM, which rely on assumptions that hold only on the traditional architectures. We provide further evidence on this in Section 5.
Figure 8. Number of programs where each optimization pass exhibits: (1) a performance gain on x86 but a loss on RISC Zero, (2) a gain on both x86 and RISC Zero with ¿5% higher gain on x86, (3) a gain on both with ¿5% higher gain on RISC Zero, or (4) a gain on RISC Zero but a loss on x86.
Figure 8 illustrates passes with divergent effects on x86 vs. RISC Zero execution. While most passes benefit both platforms, the performance improvement is generally more pronounced on x86. This exemplified by passes such as inline, simplifycfg, and jump-threading. Some passes, such as reg2mem and loop-extract, tend to benefit x86 but degrade performance on RISC Zero. In contrast, passes like ipsccp and attributor show stronger positive effects on RISC Zero. The corresponding plots for SP1 and proving time follow similar trends and are omitted for brevity.
5. Performance Impact Analysis
Based on our study’s results (Section 4), we examine a selection of passes with interesting performance behaviors. We select these passes based on four criteria: (1) passes that cause considerable performance losses on zkVMs, (2) passes that provide significant performance gains on zkVMs, (3) passes that benefit x86 execution but degrade performance on zkVMs, and (4) passes that improve both x86 and zkVM performance, but with a substantially greater impact on x86.
const N: usize = ...;
let mut v = [0; N];
for k in 0..N {
  v[k] = 42;
(a) A simple loop of depth one.
const N: usize = ...;
let mut v = ...;
for k in 0..N {
 for j in 0..N {
  for i in 0..N {
   for l in 0..N {
    v[k][j][i][l] = 42;
(b) Nested loops of depth four.
Figure 9. Simplified examples that cause significant performance regressions when using licm.
[…]
Figure 9(a) illustrates a simplified loop extracted from npb-lu where licm incurs this overhead. This comes from the increased number of getelementptr instructions in the LLVM IR, each computing the address of an array element to store the constant 42 (line 5). These extra memory operations in turn cause significant paging pressure with an increase in page-ins and page-outs. Nested loops exacerbate this effect: with a nesting of depth 4 (Figure 9(b)), licm increases cycle count by $2.6\times$, compared to $1.3\times$ for depth 2. Overall, the extra memory pressure stemming from licm is evident when loops iterate over arrays.
Insight 1 (I1): Optimizations that introduce a lot of paging pressure should be avoided. Optimizations that increase paging pressure harm zkVMs. Examples of passes that can potentially increase paging pressure include licm, inline (see our analysis below), and reg2mem. Since page-ins and page-outs are highly expensive in zkVMs, compilers targeting zkVMs must prioritize paging-aware cost models when transforming loops or spilling to memory.
pub fn work(x: u64) -> u64 {
  let mut sum = x;
  for j in 0..100 {
    sum = sum.wrapping_mul(31).wrapping_add(j);
  sum
fn main() {
  let n = 1000;
  for i in 0..n {
    let res = work(i as u64);
Figure 10. Example Rust function that inline causes a performance regression for zkVMs.
Significant improvements on zkVMs. According to RQ1 (Section 4.1), inline is the most beneficial pass for zkVMs. But is LLVM’s default inlining optimal (Theodoridis et al., 2022)? To find out, we raise the inlining threshold from 225 (default) to 4328 using the -inline-threshold option, allowing the compiler to inline more aggressively even larger functions. Notably, the value of 4328 comes from the OpenTuner’s autotuning results on the NPB and crypto benchmarks (Section 4.2). Compared to -O3 with the default threshold, the higher -inline-threshold improves execution by 6% on RISC Zero and 1% on SP1 on average, with some benchmarks showing dramatic gains, such as +44% on RISC Zero and +40% on SP1 for the npb-bt program. In contrast, increasing the inline threshold degrades x86 performance on average by 1%.
pub fn matmul(mat: &[[f64; 5]; 5], vec: &[f64; 5]) {
  let mut res = [0.0; 5];
  for col in 0..5 {
    for row in 0..5 {
      res[row] += mat[col][row] * vec[col];
Figure 11. Example Rust function that computes the matrix-vector product of a $5\times 5$ matrix and a 55-dimensional vector.
Insight 2 (I2): Increasing the inlining threshold on zkVMs can be very beneficial. A more aggressive inlining policy can yield substantial performance gains on zkVMs because it may reduce call overhead and paging costs. This suggests that compilers targeting zkVMs should: (1) adopt a zkVM-aware cost model for inlining rather than reusing heuristics from traditional CPUs, and (2) tune inlining thresholds per workload, especially for performance-critical code (Section 6.2).
The importance of tuning inlining and using a zkVM-aware model is exemplified by the program in Figure 10 (extracted from the tailcall program). Here, inlining causes major regressions on both zkVM execution ($0.8\times$ speedup) and proving ($0.45\times$ speedup). The root cause is stack spilling: RISC-V has only 32-bit registers, so each 64-bit variable occupies two. After inlining the function work (lines 1 and 12), three u64 variables (outer-loop index i on line 11, inner-loop index j on line 3, and sum on line 2) must coexist. This forces the compiler to spill both j and sum to the stack every inner-loop iteration. This doubles the lw/sw (load/store) instructions, which account for most of the $2\times$ cycle increase. Following I1 (Insight 1), inlining should be avoided when it triggers stack spills, as loads and stores may worsen execution due to expensive page-ins and page-outs in zkVMs (ris, 2025a).
Improvements on x86; degradations on zkVMs. On polybench-durbin, the loop-unroll pass improves x86 execution by 9% but degrades RISC Zero performance by 8.7%. Figure 11 shows an oversimplified version of this benchmark where loop-unroll produces this divergence between x86 and zkVM performance. Overall, x86 sees gains of 9% when loop unrolling this example program compared to the unoptimized baseline, whereas both zkVMs consistently suffer slowdowns of 3%–10% in both execution and proving time.
Table 2. The relative speedup percentage of 4x and 16x unrolling compared to the non-unrolled assembly code.
[…]
4 28.10% 24.25% 6.60% 51.43% 39.51%
16 31.51% 26.42% 1.54% 51.62% 52.66%
When used with -O3 and a high unroll-threshold (aggressive unrolling), loop unrolling becomes beneficial compared with -O3 and no loop unrolling. Specifically, on RISC Zero, execution improves by 16% and proving by 3.7%; on SP1, by 0.5% and 1.7%. However, these gains are still modest compared to x86’s $32\times$ speedup improvement.
To isolate the effect, we manually unroll loops in RISC-V assembly, avoiding interference from other optimizations or vectorization. Table 2 reports speedups for $4\times$ and $16\times$ loop unrolls. This manual effort yields substantial performance gains for both zkVMs, with RISC Zero even surpassing the gains seen on x86. These improvements come mainly from executing fewer loop bookkeeping instructions, such as counter increments, comparisons, and branch jumps.
Insight 3 (I3): Performance gains are observed only when loop unrolling leads to a reduction in cycle count. Since zkVMs cannot exploit instruction-level parallelism, out-of-order, or superscalar execution, performance gains when unrolling loops depend entirely on reducing instruction count. This suggests that compilers targeting zkVMs should adopt a tailored cost model for loop unrolling, prioritizing instruction reduction over traditional hardware assumptions.
fn abs_i32_branchy(x: i32) -> i32 {
  if x < 0 { -x } else { x }
(a) Rust function computing $|x|$ using a branch.
bltz a0, .LBB52_2
ret
.LBB52_2:
neg a0, a0
ret
(b) RISC-V before applying simplifycfg.
srai a1, a0, 31
xor a0, a0, a1
sub a0, a0, a1
ret
(c) RISC-V after applying simplifycfg.
Figure 12. An example Rust function including its assembly output before and after applying simplifycfg.
Substantial improvement on x86; minimal effect on zkVMs. On polybench-nussinov, simplifycfg improves x86 by 23.6%, but SP1 execution by only 2.6%. Figure 12 shows a minimized function from this benchmark that computes the absolute value of an integer using a branch. simplifycfg replaces the branch with equivalent arithmetic operations (Figures 12(b) and 12(c)). When tested with an array of integers chosen at random, the optimized version runs $2.2\times$ faster on x86, presumably thanks to less branch misprediction overhead. However, on zkVMs, the optimized code increases cycle count by 17.7% on RISC Zero and 7.6% on SP1. Proving time regresses similarly (18.1% and 7.0%), and execution time degrades by 8.5% and 7.7%, respectively.
Insight 4 (I4): branches should not always be eliminated. Compilers targeting zkVMs should be more conservative when removing branches, for two reasons. First, branches are relatively cheap and incur no misprediction penalty compared to traditional architectures. Second, predicated execution increases proving cost, since both paths must be proven.
6. Implications
We demonstrate the practical value of our study’s results by refining existing LLVM passes to improve zkVM performance and then distill several key lessons from our findings.
6.1. zkVM-Specific Optimizations
Existing LLVM optimization levels already boost zkVM performance (Section 4.2), but our root cause analysis (Section 5) reveals significant untapped potential. Building on these insights, we propose a non-exhaustive set of lightweight backend and pass modifications across three dimensions to further unlock LLVM’s potential for zkVMs.
Change set 1: Adopting a zkVM-specific cost model. The default RISC-V cost model in LLVM is not zkVM-aware, yet instruction costs on zkVMs differ greatly from traditional CPUs. For example, multiplication costs the same as addition, division is not expensive, while floating-point operations are extremely costly due to emulation (ris, 2025b). Memory access is also far more expensive when it triggers page-ins or page-outs (see Insight 1, Section 5). We address these by updating the RISCVTTIImpl and RISCVTargetLowering classes, which provide target-specific cost models to LLVM’s middle-end so that passes can make better decisions for this architecture. Our modifications encode a zkVM-aware cost model for arithmetic and memory operations (e.g., division is no longer considered expensive).
Change set 2: Updating defaults and heuristics. Many LLVM passes use default parameters and heuristics that are tuned for traditional architectures. This tuning may not be suitable for zkVMs. Adjusting these for zkVM-specific instruction costs, such as making inling agressive (Insight 2) or applying loop unrolling only when it reduces instructions count (Insight 3), can improve generated code performance.
We increase the default inlining threshold and adjust parameters such as inline-call-penalty accordingly. We also update simplifycfg to make branch elimination more conservative and avoiding evaluation of both sides of a predicated operation (Insight 4).
Change set 3: Disabling irrelevant passes. Our RQ1 analysis (Section 4.1) shows that some passes offer little or no benefit on zkVMs, mainly because they rely on hardware features that are unsupported in zkVMs. For example, speculative-execution hoists side-effect-free instructions to reduce branch-misprediction latency on out-of-order, superscalar CPUs. This advantage is irrelevant to zkVMs, which execute RISC-V code strictly in order inside an arithmetic circuit. Similarly, we also disable loop-data-prefetch, and hot-cold-splitting, as they provide no measurable gain on zkVMs.
All these targeted LLVM changes took only a few hours to implement and required fewer than 100 lines of code.
Figure 13. Changes in zkVM execution performance -O3 with our modified LLVM compared to -O3 with the original LLVM. Programs with effects below 2% are omitted.
[…]
The largest regression occurs when proving regex-match on SP1 (+27.3%). This is caused by our changes in the inline’s heuristics, which make SP1 produce 20 proof shards instead of 16. Proof sharding divides the program into smaller proving units; more shards increase proving time due to additional aggregation overhead.
6.2. Lessons Learned and Discussion
zkVM users can rely on the compiler to optimize their programs. Our results show that zkVM users can already benefit from LLVM’s existing optimization levels (e.g., -O3), which improve both execution and proving time by over 40% (RQ2, Figure 5). However…
…the magnitude of these gains is smaller compared to what similar optimizations achieve on traditional architectures. Our fine-grained pass-level analysis (Section 4.1) reveals that many passes still rely on hardware-centric assumptions and heuristics that harm zkVM performance (e.g., branch elimination, Section 5). By refining just a few of these passes to be zkVM-aware, we already surpass LLVM’s default -O3 (Figure 13). This proof-of-concept highlights that even small, yet targeted zkVM-specific compiler changes can deliver substantial speedups and motivate the development of a dedicated zkVM-aware LLVM backend.
Section 5 covers only a few passes (e.g., licm, inline). Many others could unlock further optimization opportunities. For example, instcombine improves half the benchmarks but hurts the rest (Figure 4). Performing a similar root-cause analysis on this pass could reveal which rewrites of instcombine to keep and which to disable.
Autotuning performance-critical software. As part of RQ2 (Section 4.2), we also show the viability of program autotuning on zkVMs. This has direct, tangible value in real-world settings. For example, in the Ethproofs ecosystem (Ethereum Foundation, 2025), where different zkVMs race to prove Ethereum blocks in real-time (i.e., in less than 12-seconds – Ethereum’s block time). Despite significant progress, real-time proving still requires substantial GPUs, and much research and engineering work focuses on achieving even minor improvements. Exploring autotuning for prover workloads on real Ethereum blocks is thus a promising and economically impactful direction.
Another takeaway from our autotuning experiment is that it can yield significant gains even for programs using precompiles, with improvements of up to 10%. This suggests that compiler-level tuning can complement zkVMs’ built-in accelerations by further reducing proving costs.
Profile-guided optimization. High-quality profiling data is critical for generating efficient code. For example, static profile-guided optimization could significantly improve inlining decisions and reduce instruction count. Profiling data could also guide the design of future zkVM-specific passes. For example, a page coalescing pass could place hot variables within as few 1 KB pages as possible, minimizing expensive page-ins and page-outs on zkVMs, which we identify in Section 5 as a major bottleneck.
Use of superoptimization. Since zkVM execution and proving are orders of magnitude more expensive than native execution, it is often worth trading longer compilation time for faster proving. Superoptimization is a technique that allows compilers to do so by finding the optimal sequence of instructions for implementing a given computation. Superoptimizers that operate on LLVM IR (e.g., Souper (Sasnauskas et al., 2018)) could be adapted to zkVMs. Incorporating a zkVM-specific cost model would make them a powerful tool for generating highly efficient zkVM guest code.
Optimized code may uncover bugs in zkVMs. In RQ2 (Section 4.2), autotuning a specific program exposed SP1 bugs, including a security-critical case where the zkVM silently aborted mid-execution yet still produced a proof that passed verification. Such issues show that optimized code can stress parts of a zkVM that unoptimized inputs miss. These bugs highlight the need to improve zkVM reliability, which is also exemplified by bug bounty programs for many zkVMs (HackenProof, Inc., 2025; code4arena, 2025). Inspired by compiler testing (Yang et al., 2011; Le et al., 2014), zkVM testing could use optimized vs. unoptimized runs as a test oracle, that is, two equivalent versions should produce identical results, with any divergence indicating a bug.
7. Related Work
zkVMs. Ben-Sasson et al. (2014) introduce the first system to prove executions of a von Neumann machine, albeit on a bounded machine with a predetermined cycle count. The field has since moved to unbounded proving, e.g., Cairo (Goldberg et al., 2021) which leverages STARKs (Ben-Sasson et al., 2018) to scale. Recent efforts improve zkVM performance via better circuit generation and prover back-ends as well as more efficient proof systems (Arun et al., 2023; Setty et al., 2023; Goldberg et al., 2021; Labs, 2024b; Bruestle and Gafni, 2023). Our work complements these improvements by studying how compiler optimizations interact with zkVM cost models, and outline directions to reduce proving overhead through traditional (or novel) optimizations.
Benchmarking ZKPs. Benchmarking ZKPs is crucial for identifying bottlenecks and guiding improvements. Ernstberger et al. (2024a) introduce zk-bench, a framework for benchmarking ZK DSLs (Bellés-Muñoz et al., 2023). Chaliasos et al. (2024b) provide the first large-scale systematic study of ZK-EVMs, highlighting design trade-offs and bottlenecks. Industry efforts benchmark zkVMs to compare capabilities and performance (a16z crypto, 2024; Labs, 2024d; Inc., 2024). We build on these benchmarks by reusing selected industry workloads and complement them with a broad and diverse Rust/C suite.
Autotuning & Superoptimization. Ren et al. (2021) develops BinTuner to investigate the impact of compiler optimizations on binary code difference. BinTuner is based on OpenTuner (Ansel et al., 2014), a framework for building domain-specific autotuners using genetic algorithms. In this work, we leverage OpenTuner to show that autotuning is a viable method to improve the proving performance of zkVMs for a given program. A complementary line of work is superoptimization (Schkufza et al., 2012; Sasnauskas et al., 2018; Massalin, 1987), which searches for optimal instruction sequences for a given (typically loop-free) fragment. Exploring superoptimization for zkVMs is a promising future work, and our study provides valuable insights to guide it (Section 6.2).
8. Conclusion
We presented the first systematic study of how compiler optimizations affect two production RISC-V zkVMs. Most passes help, yet gains are modest and a few are harmful (e.g., licm). We established a strong correlation between cycle count and proving time, enabling cycle-count-guided autotuning that yields up to $2.2\times$ proving speedups beyond -O3, and we prototyped zkVM-aware LLVM changes that further improve the performance of zkVMs.
Future work can extend this study to additional zkVMs and proof systems, unifying cost metrics (e.g., prover-gas-like models), and exploring superoptimizations for zkVMs. These compiler-centric directions can push zkVM proving performance beyond what proof-system research and engineering alone can deliver.
Appendix A Differences between zkVMs and Native Execution
The following considers two popular zkVMs (RISC Zero (Bruestle and Gafni, 2023) and SP1 (Labs, 2024b)) and summarizes the key differences compared to traditional architectures.
Instruction latency. On traditional architectures, instruction latency can vary significantly depending on the instruction type, cache hits/misses, and other micro-architectural effects (Fog, 2022). In zkVMs, however, most instructions have a uniform cost (ris, 2025a), with most instructions taking exactly the same number of cycles.
Memory access. Memory access is relatively cheaper on zkVMs. For example, on RISC Zero, memory access costs only one cycle if the page is already paged in. Page-ins and page-outs on RISC Zero are, however, more expensive, taking around 1130 cycles on average (ris, 2025a). zkVMs also do not have any caches making memory access always synchronous.
Control flow. zkVMs do not have branch prediction or instruction caches, which means that control flow instructions can be cheaper on zkVMs, as there is no penalty for mispredicted branches.
Out-of-order execution. zkVMs do not have out-of-order execution. They also do not have super-scalar execution, instruction-level parallelism, or pipelining, which means that all instructions are executed sequentially.
Floating-point operations. zkVMs do not support native floating-point operations, which means that all floating-point operations must be emulated. This can lead to significant overhead, especially for complex floating-point operations.
Multithreading. zkVM execution is always single-threaded. There is no parallelism in the execution.
Precompiles. zkVMs provide precompiles for expensive operations such as cryptographic hashes or elliptic-curve arithmetic. Precompiles can significantly reduce the cost of these operations compared to implementing them in the guest program.
Figure 14. Median duration of zkVM operations compared to execution time on x86 for the NPB benchmark suite. All binaries were compiled with no optimizations.
zkVM execution and proving are order of magnitude slower than native execution. Figure 14 illustrates this by considering the time spent on zkVM execution, proof generation, and native execution for every program included in the NPB benchmark suite (Martins et al., 2025).
Table 3. Benchmark programs and their SLOC counts, split across two subtables with balanced rows.
Program Libraries used Precomp. Rust Total
bigmem None No 32 32
ecdsa-verify k256 Yes 16 16
eddsa-verify ed25519_dalek Yes 16 16
factorial None No 26 26
keccak256 sha3 Yes 13 13
loop-sum None No 29 29
merkle rs_merkle No 46 46
npb-bt None No 2915 2915
npb-cg None No 827 827
npb-ep None No 485 485
npb-ft None No 1151 1151
npb-is None No 815 815
npb-lu None No 2746 2746
npb-mg None No 2121 2121
npb-sp None No 2728 2728
polybench-2mm None No 113 113
polybench-3mm None No 148 148
polybench-adi None No 107 107
polybench-atax None No 78 78
polybench-bicg None No 80 80
polybench-cholesky None No 63 63
polybench-correlation None No 98 98
polybench-covariance None No 84 84
polybench-deriche None No 146 146
polybench-doitgen None No 83 83
polybench-durbin None No 67 67
polybench-fdtd-2d None No 93 93
polybench-floyd-warshall None No 59 59
polybench-gemm None No 90 90
polybench-gemver None No 116 116
Program Libraries used Precomp. Rust C/C++ Total
polybench-gesummv None No 81 81
polybench-gramschmidt None No 88 0 88
polybench-heat-3d None No 82 0 82
polybench-jacobi-1d None No 64 0 64
polybench-jacobi-2d None No 70 0 70
polybench-lu None No 64 0 64
polybench-ludcmp None No 103 0 103
polybench-mvt None No 79 0 79
polybench-nussinov None No 99 0 99
polybench-seidel-2d None No 70 0 70
polybench-symm None No 88 0 88
polybench-syr2k None No 85 0 85
polybench-syrk None No 81 0 81
polybench-trisolv None No 67 0 67
polybench-trmm None No 74 0 74
regex-match regex No 45 0 45
rsp serde, Yes 11 0 11
rsp-client-executor,
c-kzg,
bytemuck_derive,
bincode
sha2-bench sha2 No 36 0 36
sha2-chain sha2 No 44 0 44
sha3-bench sha3 No 35 0 35
sha3-chain sha3 No 44 0 44
sha256 None No 163 0 163
spec-605 None No 48 2159 2207
spec-619 None No 53 203 256
spec-631 None No 46 6496 6542
tailcall None No 56 0 56
zkvm-mnist None No 181 0 181
Appendix B Benchmark Suite Details
In the following, we give a brief summary of the benchmark programs.
PolyBench. PolyBench is a benchmark suite of 30 numerical computations from domains like linear algebra, image processing, physics simulations, dynamic programming, statistics, and more. Originally written in C (Pouchet, 2010), we use the Rust port (Ferrer, 2020) of the benchmark suite. The inputs of the benchmarks have been reduced to fit within the constraints of the zkVMs.
NPB. The NAS Parallel Benchmarks (NPB) are a set of eight benchmarks designed to evaluate the performance of parallel supercomputers originally developed by NASA’s Advanced Supercomputing Division (Division, 2024). Martins et al. (Martins et al., 2025) port the benchmarks to Rust and provide a sequential version of the benchmarks. We use the sequential version of the Rust port for our evaluation. Similar to PolyBench, the input sizes have been reduced to fit within the constraints of the zkVMs.
SPEC CPU 2017. The SPEC CPU benchmark suite (Bucek et al., 2018) is a widely used benchmark suite for evaluating the performance of computer systems. Due to the complexity of the benchmarks and the fact that they are written in C/C++, whose support is still limited in zkVMs, we only use a subset of three benchmarks: 605, 619 and 631. These three benchmarks were chosen based on their ease of integration.
a16z crypto zkVM benchmarks. Andreessen Horowitz (a16z) crypto has published a set of benchmarks (a16z crypto, 2024) comparing the RISC Zero, SP1 and Jolt (Arun et al., 2023) zkVMs. We use the benchmarks from their repository, which consist of a set of cryptographic operations (sha2, sha3) as well as an allocation-heavy benchmark (bigmem).
Succinct Labs benchmarks. We also incorporate a subset of the zkVM benchmarks released by Succinct Labs (Labs, 2024d). The benchmarks we include consist of a set of cryptographic operations (ecdsa-verify, eddsa-verify, keccak256) as well as a Fibonacci sequence benchmark (fibonacci).
Reth Succinct Processor (RSP). We also adopt the Reth Succinct Processor (RSP) (Labs, 2024a) benchmarks from Aligned Inc.’s benchmark suite (Inc., 2024). RSP represents a common type of workload for zkVMs that generates zero-knowledge proofs for EVM (Ethereum Virtual Machine) block execution. We use block 20526624 as input for our benchmarks. [^1] The FAQ on RSP’s GitHub page highlights block 20526624 as ”a good small block to test”. While adapting our benchmark setup to test a different block is relatively straightforward, running the benchmarks for all optimization profiles is time-consuming. The impact of compiler optimizations on RSP under consideration of different blocks is left as future work.
Others. In addition to the above benchmarks, we also add some additional benchmarks: a benchmark training a neural network on the MNIST dataset (zkvm-mnist) downsampled to 7x7 images, a benchmark performing regex matching (regex-match), a benchmark performing a merkle tree inclusion proof (merkle), another SHA-256 benchmark (sha256) as well as three smaller benchmarks (factorial, loop-sum and tailcall).
Overall, the 58 benchmark programs that we have assembled span a diverse range of workloads and input characteristics, ranging from numerical computations to memory-intensive operations, cryptographic primitives commonly used in zkVMs, and real-world applications such as RSP. This variety ensures a thorough evaluation of compiler optimizations across all major zkVM use cases.
Benchmark Programs. Table 3 summarizes the benchmark programs used in this evaluation, including their source lines of code (SLOC) counts for Rust and C/C++. The SLOC counts of any libraries used by the benchmarks are not included in the table.
A list of the library versions is listed in Table 4. All programs also use the crates sp1-zkvm (version 4.1.1) and risc0-zkvm (version 1.2.3).
Table 4. Programs and their associated libraries and versions.
Program Library Version/Commit
PolyBench (Ferrer, 2020) babdd97
NPB (Martins et al., 2025) 4bd879b
ecdsa-verify k256 0.13.3
eddsa-verify ed25519_dalek 2.1.1
merkle rs_merkle 1.5.0
regex-match regex $1.1$
sha2-bench sha2 0.10.8
sha2-chain sha2 0.10.8
Program Library Version/Commit
sha3-bench sha3 0.10.8
sha3-chain sha3 0.10.8
rsp serde 1.0.204
rsp-client-executor 249b34e (SP1)
rsp-client-executor 4ceefdf (RISC Zero)
c-kzg 1.0.3
bytemuck_derive 1.8.0
bincode 1.3.3
Appendix C Statistics of Baseline
Table 5 shows the statistics of execution and proving times (in seconds) for each zkVM across all benchmarks, measured on the unoptimized baseline with no compiler passes applied.
Table 5. Statistics of execution and proving times (in seconds) for each zkVM across all benchmarks, measured on the unoptimized baseline with no compiler passes applied.
zkVM Execution Proving
Min Max Mean Median Min Max Mean Median
RISC Zero 0.04 157.70 4.51 0.34 0.53 2071.24 60.85 3.83
SP1 0.06 41.81 1.70 0.23 0.38 205.87 8.89 1.90